# Decentralization of Ethereum’s Builder Market

### v3

@@ Abstract
The builder market was introduced to fairly distribute Maximal Extractable Values (MEV) among validators and avoid validator centralization. As of the time of writing, three builders produced the vast majority (more than 80%) of blocks in Ethereum, creating a concerning centralization factor. However, the community believes that such centralization “is okay,” arguing that builder centralization will not lead to validator centralization. In this empirical study, we interrogate the causes and implications of builder centralization and challenge this belief that it is acceptable.
Our main finding is that builder centralization has led to a significant loss by validators and, if left uncontrolled, could lead to validator centralization. Moreover, MEV mitigation solutions slated for adoption are affected too because they rely on the builder market as an “MEV oracle,” which is made inaccurate by centralization.
Our investigation revealed two reasons behind builder centralization. We propose a structural change to the existing MEV supply chain and a solution based on the new supply chain structure. However, future work is required to analyze if the new supply chain structure is sustainable in the long term, which we leave open.
@@ 1 Introduction
At the core of any decentralized blockchain, a key assumption underpinning their strong security properties (integrity, availability, censorship resistance, incentive compatibility, etc.) is “decentralization”, that the system is run by a large number of independent participants [63, 28, 75, 83, 4, 74]. It is community consensus that one of the greatest threats to decentralization is Maximal Extractable Value (MEV) [16], profits that protocol participants, known as validators, can reap by manipulating the ordering of transactions. MEV can lead to centralization because it disproportionally benefits well-resourced validators compared to “regular” ones, as extracting MEV requires significant capital, intelligence, and computational resources, which regular validators may not have [4, 23, 67, 12, 11].
@@
Figure 1: Market share of builders. The builder market is one of the least decentralized parts of Ethereum!
@@
Is the builder market working as intended? The answer is rather unclear because the builder market is highly centralized. As shown in Fig. 1, three builders won over 90% of auctions, raising questions about the competitiveness of the market. However, the community believes that “it is OK to centralize block building” [23], arguing that block centralization will not lead to validator centralization. We challenge this belief through an empirical study of the past MEV-Boost auctions. As will be shown in detail, builder centralization has caused a significant loss to the proposers and, if left uncontrolled, could lead to undesired consequences including validator centralization.
1.1 Implications of Builder Centralization
@@
Proposer loss and its consequences.   Intuitively, a proposer’s loss is the difference between the MEV available in a block and the auction revenue. Builder centralization can cause proposer loss in two ways. First, auctions may not be competitive enough due to, e.g., collusion or other limitations of the auction mechanism. Second, even if builders compete with each other, if a single builder has a unique advantage, it will not be incentivized to bid on all of the MEV it can extract. Instead, it only needs to outbid the second-best bidder slightly to win. We say that the proposer incurs a loss from uncompetitiveness in the first case and a loss from inequality in the second case.
@@
If proposers suffer significant losses, they would be incentivized to extract MEV by themselves (instead of using the builder market). This is undesirable because it re-introduces the dynamics that PBS wants to avoid: sophisticated validators would have an advantage over small ones in extracting MEV. In other words, significant proposer losses would make PBS “unstable”. Moreover, several MEV mitigation solutions (e.g., MEV-burn [19]) rely on the MEV-Boost auctions as an “oracle” to know how much MEV has been extracted from the system (e.g., MEV-burn proposed to remove that amount from circulation, effectively redistributing them to all holders of ETH through deflation.) The proposer’s loss translates to inaccuracy in the MEV oracle.
@@
Next, we introduce the auction mechanism and present our method to quantify the proposer losses using auction data.
@@
Modeling MEV-Boost auctions.   MEV-Boost auction adopts an open-bid, ascending price auction akin to an English auction [6], with a 12-second deadline. For each auction instance, builders submit bids in the form of $(B,BV)$ specifying a block $B$ and a bid value (BV) the builder is willing to pay for $B$ to be added to the blockchain. The revenue a bidder gains from $B$ is called the true value ($TV(B)$), which can be estimated from transactions in $B$. At the end of an auction, the block $B$ with the highest bid is added to the blockchain, the builder pays $BV$ to the proposer and earns $TV$ when $B$ is executed (i.e., the builder’s profit is $TV-BV$).
@@
Recall that the proposer’s loss is the difference between the auction revenue and the MEV available in a block. The latter can be represented by the builders’ highest true values. With that, the two types of losses can be computed from auction data, as illustrated in Fig. 2. Specifically, suppose the true values of the bids from a given auction instance are $TV_{1}>TV_{2}\geq\cdots\geq TV_{N}$. The loss from inequality is $TV_{2}-TV_{1}$ because a competitive auction can, at most, generate revenue of $TV_{2}$. The loss from uncompetitiveness is $\max(TV_{2}-BV,0)$ (if this value is negative, the auction is considered competitive, and there is no loss from uncompetitiveness.)
@@
1.2 What caused Inequality in Block-building Capacities
@@
In the current MEV ecosystem (to be presented in Sec. 2.2), builders receive order flows from order flow providers and compute the optimal ordering of received order flows (see Appx. E for a reference builder algorithm). Thus, a builder’s block-building capacity is determined by the “quality” of the order flows it receives. At the same time, not all order flows are accessible to all builders. The difference in builder’s ability to access important private order flows drives the inequality in block-building capacity. This also explains builder centralization: only a small set of builders can access these profitable order flows.
We proposed a metric to identify important order flows, and we investigated their accessibility by different builders and the underlying reasons.
Identifying pivotal private order flow providers.   To formally capture the intuition of “important providers”, we defined pivotal level to quantify the sustained impact of a given provider. A provider $P$ is pivotal in an auction $A$ if the winner of $A$ would have lost, had transactions from $P$ were not available. In other words, pivotal providers are necessary for the winner to win. The pivotal level of a provider $P$ is the percentage of MEV-Boost auctions in which $P$ is pivotal.
@@
These order flows are not equally inaccessible to all builders for two reasons we now introduce.
Inequality due to trust barriers.   Order flow providers do not openly share private order flows with any builder because a malicious builder can harm the provider by, e.g., sandwich attacks [94] and imitation attacks [71]. Providers set up different access barriers for self-protection.
Most MEV searchers (a specific type of order flow providers) resort to “obscurity for security” in that they are not publicly reachable; thus, we could not learn about their access barriers and how to overcome them. On the other hand, MEV channels (another type of provider) typically publish a clear access policy, which allowed us to gain some insight into the nature of their access barrier.
For example, MEV-Blocker only shares order flows with builders with a market share of at least 1%, using market share as a proxy for reputation and trustworthiness. This requirement introduces a chicken-and-egg problem for new builders: to gain market share, they need access to private order flows to win auctions, but as newcomers, they lack the required market share to access these flows. As a result, new builders often resort to subsidization: bidding higher than their actual profits in MEV-Boost auctions to establish the required market shares. Therefore, a new builder needs to keep subsidizing until reaching a certain market share to overcome the access barriers set by MEV channels. The cost to do so continues to increase over time and can reach up to 7.4 ETH, according to our quantification.
In summary, with a moderate cost, builders can overcome the access barrier of one type of order flow provider (MEV channels). However, how to access order flows from other providers, particularly important searchers, remains unclear.
Inequality due to integration.   A more worrisome finding is that some providers send order flows almost exclusively to their preferred builders, causing an inequality in block-building capacities.
To identify such provider-builder pairs, we introduced builder-specific pivotal level of a provider $P$ for a builder $B$ as the pivotal level of $P$ in auctions won by $B$. With this metric, we rediscovered known exclusive relationships (called “integration” [40]), but also found unreported integrated providers for top builders (Titan and Beaver respectively). This finding is interesting because Titan has been the only top builder without a well-known integrated provider. Our findings show that all three top builders, accounting for 90% of market share, have integrated providers.
@@
1.3 Mitigation Ideas
To summarize, we found that trust barriers and incentives for integration contribute to builder centralization, leading proposers to incur significant losses. The consequences include the instability of PBS and the inaccuracy of MEV oracles.
We propose a structural change to the existing MEV supply chain (Sec. 2.2) and a solution based on the new supply chain structure. However, future work is required to analyze if the new supply chain structure is sustainable in the long term, which we leave open.
The new supply chain has only one builder to which the order flow providers send flows and from which the proposers accept blocks. In this setting, a recent work [57] suggests a truthful mechanism that incentivizes order flow providers to report the MEV in their order flows truthfully.
Our solution involves having the builder run the truthful mechanism in [57] in a Trusted Execution Environment (TEE) [15], hardware-based enclaves that guarantee the confidentiality and integrity of programs running inside. Order flow providers verify that the correct mechanism is running in the TEE using remote attestation, then send order flows to the TEE along with a bid indicating the MEV in their order flow. Then, the TEE computes a fair refund to order flow providers. The refund mechanism is designed so that bidding truthfully is a dominant strategy [57]. (While oversimplifying, one can make a loose analogy to sealed-bid second-price auctions among order flow providers.) Then, the TEE can compute the total MEV in a block from the providers’ bids, solving the problem of inaccurate MEV oracles.
This also addresses the instability of PBS because the builder program in the TEE can be programmed to pay the proposer fairly, leaving just enough for its operation. Since this builder is fair to order flow providers, it is conceivable that it will be preferred by order flow providers over other builders. However, as mentioned above, it is unclear if having a single builder is an equilibrium. Also, the proposers may be incentivized to become order flow providers, which may re-introduce the concern similar to PBS instability. We elaborate on the idea in Sec. 7.1.
Contribution
To the best of our knowledge, we are the first to use the rich auction data to analyze the decentralization of Ethereum’s builder market, while most previous works are limited to on-chain data. We summarize our contribution below.
  • We found that builder centralization has led to significant proposer losses, which could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles. We challenge the community’s belief that “builder centralization is okay".
@@ 1.3 Mitigation Ideas
  • We identify two reasons for inequality in block-building capacity: providers’ trust concerns and the incentive for integration.
  • Our dataset of the true values of 301,479 MEV-Boost auctions from April to December 2023 may be of independent interest. Our dataset of 8.79 billion partial bids since 2022 is the largest scale known to us.
  • Besides contributing to the literature on MEV mitigation, our work also contributes to the Economics literature on the analysis of MEV-Boost auctions. Currently, there is only limited auction theoretic analysis that considers the complexity of real-world MEV auctions [40, 67, 90]. The findings in this paper can serve as empirical evidence for future theoretical interrogation.
@@ 2.1 MEV-Boost Auctions
PoS Ethereum.   Before introducing background on MEV-Boost auctions, we first cover the relevant concepts in Ethereum’s Proof of Stake-based consensus protocol called Gasper [13]. Participants of Gasper are called validators. Each validator must deposit 32 ETH as collateral [22]. Time is divided into 12-second intervals called slots. This results in 7,200 slots in Ethereum per day. During each slot, a single validator, referred to as the proposer, proposes a block, while other validators vote for the chain head in their local view.
@@ 2.2 MEV Supply Chain
Channel.   A private channel, or channel for short, connects users directly with builders, bypassing the mempool. Channels receive transactions from users and then, forward them to the user-specified builders, offering privacy and atomicity guarantees. E.g., a user can send a bundle (list) of transactions and request them not to be unbundled. Examples include MEV-Share [37], MEV Blocker [60], BloXroute [9], etc.
Searcher.   Searchers are a special type of users. They run (typically proprietary) algorithms to construct profitable transaction sequences (bundles) based on the pending transactions and the blockchain state. Same as regular users, searchers may submit transactions to mempool or channels. In addition, most builders have public APIs that allow searchers to submit bundles directly. Some searchers may prefer channels for the privacy and atomicity benefits. A searcher is said to integrate with a builder if it sends almost all of its orders to that builder exclusively. For example, searcher 0xA69b… is considered an integrated searcher of Beaver Builder because Beaver Builder includes the majority of its transactions.
@@
Note that $EV(B)$ is a lower bound of the builder’s actual income from $B$ because the builder may own accounts out of $A_{p}$ in Equation 1, or even receive payments off-chain. We discuss the implications in Sec. 7.
@@ 3 Dataset
We created two auction datasets to study the decentralization of the builder market: a dataset of partial bids from September 2022 to July 2024, and a dataset of full bids from April to December 2023. The full bids are obtained from the ultra sound relay [62]. Recall from Sec. 2.1 that a bid consists of a block $B$ and a bid value $BV$ and a partial bid consists of $(\mathsf{Hash}(B),BV,PK_{\text{builder}},metadata)$ where $metadata$ includes parent hash, gas limit, gas used, etc. To support analysis, we also collected on-chain data, including blocks and transactions on Ethereum, as well as miscellaneous data, which will be detailed below. Tab. 1 summarizes the datasets.
@@
On-chain blocks 4,892,579 Reth node
transactions 748,859,795
Auction blocks produced by MEV-Boost 4,305,571 relay API
partial bids 8,789,557,915 relay API
full bids from ultra sound 326,574,676 ultra sound relay
Misc. builder public keys 248 see Sec. 3.1
@@
searcher addresses 2,140
private transactions 77,747,623
transaction sources 60,345,066
Table 1: Overview of collected datasets.
@@ 3.1 Data Collection and Validation
On-chain data.   We run an Ethereum execution client Reth [68] along with a consensus client Lighthouse [69] to collect blocks and transactions from September 15, 2022, to July 31, 2024. The on-chain dataset includes 4,892,579 blocks and 748,859,795 transactions.
@@
Data from relay APIs.   We connected to all 13 known relays [^2] Three relays were inactive during the period of our study. and collected the following information from their public APIs: 1) the winning bid including the final block sent to the proposer; and 2) partial bids for blocks that did not win the auction. Recall that partial bids do not include the block body. We collected 4,305,571 winning bids and 8,789,557,915 partial bids from September 15, 2022, to July 31, 2024. We use winning bids to identify blocks produced through MEV-Boost. We use partial bids to quantify the importance of private order flow providers in Sec. 5.
@@
To validate the completeness of the data collected from relay APIs, we compared the winning bids we collected with three public datasets [64, 85, 17]. Our dataset includes 720 more winning bids than other datasets combined. Upon comparison with on-chain blocks, we find they are all included on Ethereum, thereby indicating a high level of completeness.
@@
Full bids from the ultra sound relay .   From the ultra sound relay [62], we obtained the full bids in 301,479 historical auctions that took place between April and December 2023 (ultra sound relay stopped collecting bids at the end of 2023). Since the data volume is large (we record roughly 200 GB of data per day), our dataset samples one week from each month: April 9-15 and May 1-7 to December 1-7, 2023. The dataset comprises 326,574,676 full bids (block, bid value, and builder public key). We use full bids in Sec. 4 to quantify the competitiveness and efficiency of MEV-Boost auctions.
@@
The original data from relays includes bids submitted after the deadline because of network latency. We removed them as we are only interested in the bids before the auction ends. On April 18, 2023, MEV-Boost introduced bid cancellation that allows a bidder to cancel a previous bid by submitting a lower one. However, proposers can query bids before they are canceled and store them locally [24], effectively bypassing cancellation. Therefore we do not consider cancellation.
@@
Not all builders connect to the ultra sound relay, so we need to validate that the full bids from ultra sound relay cover a significant subset of all builders. We compute the percentage of the builders that submit to ultra sound relay over all known builders. The set of all known builders is a union of all builders that appeared in our datasets, including winning bids, partial bids, and bids from the ultra sound relay. As shown in Fig. 10, bids from the ultra sound relay covered more than 80% of the builders in over 75% of the MEV-Boost auctions during our study period. Moreover, the top-5 builders by market share (Flashbots, Beaver, Rsync, Builder0x69, and Titan) almost always submitted to ultra sound relay.
@@
  • Builder public keys: We collect public keys controlled by each builder combining two data sources: the public keys in builders’ official documents; and the distinct marks in their blocks’ “extra field” (see Appx. B for details).
    To validate the completeness of our dataset, we manually cross-check it with public datasets [72, 65] and related works [42, 41], and the results show that our dataset includes 119 more public keys, so our dataset is the most complete one so far.
@@
  • Searcher addresses: We obtained known searcher addresses from Etherscan [25], libMEV [50], and related work [42].
@@
  • Transaction sources: we are interested in where a transaction originates from and which part of the MEV supply chain it has traversed through. We identified transaction sources using ground truth data (as described in Appx. B).
4 Implications of Builder Centralization
In this section, we focus on the status quo and study the implications of a centralized builder market as we have today. In particular, we are interested in the losses incurred by proposers, as significant losses could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles.
We quantify proposers’ losses in historical MEV-Boost auctions considering two factors: the competitiveness of the auctions and the inequality in block-building capacity.
Our study includes some MEV-Boost auctions where the highest bids were not included on Ethereum due to network latency. However, we do not omit these auctions since the auctions themselves remain meaningful.

### v4

@@ Abstract
The builder market was introduced to fairly distribute Maximal Extractable Value (MEV) among validators and avoid validator centralization. As of the time of writing, two builders produced more than 85% of blocks in Ethereum, creating a concerning centralization factor. However, a common belief is that such centralization “is okay,” arguing that builder centralization will not lead to validator centralization. In this empirical study, we quantify the significant proposer losses within the centralized builder market and challenge the belief that this is acceptable.
The significant proposer losses, if left uncontrolled, could undermine the goal of PBS. Moreover, MEV mitigation solutions slated for adoption are affected too because they rely on the builder market as an “MEV oracle,” which is made inaccurate by centralization. Our investigation reveals the incentive issue within the current MEV supply chain and its implications for builder centralization and proposer losses. Finally, we analyze why the proposed mitigation cannot work and highlight two properties essential for effective solutions.
@@ 1 Introduction
At the core of any decentralized blockchain, a key assumption underpinning their strong security properties (integrity, availability, censorship resistance, incentive compatibility, etc.) is “decentralization”, that the system is run by a large number of independent participants [1, 2, 3, 4, 5, 6]. It is community consensus that one of the greatest threats to decentralization is Maximal Extractable Value (MEV) [7], profits that protocol participants, known as validators, can reap by manipulating the ordering of transactions. MEV can lead to validator centralization because it disproportionally benefits well-resourced validators compared to “regular” ones, as extracting MEV requires significant capital, intelligence, and computational resources, which regular validators may not have [5, 8, 9, 10, 11].
@@
Figure 1: Market share of builders (the source of the data is introduced in Sec. 3). The builder market is one of the least decentralized parts of Ethereum!
@@
Is the builder market working as intended? The answer is rather unclear because the builder market is highly centralized. As shown in Fig. 1, two builders won over 90% of auctions. The builder centralization creates immediate security problems like censorship [4]. However, the community still believes that “it is OK to centralize block building” [8], arguing that block centralization will not lead to validator centralization. We challenge this belief through an empirical study of the past MEV-Boost auctions. As will be shown in detail, proposers have incurred significant losses in a centralized builder market and, if left uncontrolled, could lead to undesired consequences including undermining PBS.
1.1 Proposer Loss in the MEV-Boost Auctions
@@
Proposer loss and its consequences.   Intuitively, a proposer’s loss is the difference between the MEV available in a block and the auction revenue. Proposers may incur losses in a centralized builder market in two ways. First, auctions may not be competitive enough due to, e.g., collusion or other limitations of the auction mechanism. Second, even if builders compete with each other, if a single builder has a unique advantage, it will not be incentivized to bid on all of the MEV it can extract. Instead, it only needs to outbid the second-best bidder slightly to win. We say that the proposer incurs a loss from uncompetitiveness in the first case and a loss from inequality in the second case.
@@
If proposers suffer significant losses, they would be incentivized to extract MEV by themselves (instead of using the builder market). This is undesirable because it re-introduces the dynamics that PBS wants to avoid: sophisticated validators would have an advantage over small ones in extracting MEV. Moreover, several MEV mitigation solutions rely on the MEV-Boost auctions as an “oracle” to know how much MEV has been extracted from the system (e.g., MEV-burn [13] proposed to remove that amount from circulation, effectively redistributing them to all holders of ETH through deflation.) The proposer’s loss translates to inaccuracy in the MEV oracle.
@@
Given the undesired consequences of proposer losses, it is important to understand the extent of losses incurred by proposers in the current builder market. Next, we introduce the auction mechanism and present how we quantify proposer losses using auction data.
@@
Modeling MEV-Boost auctions.   MEV-Boost auction adopts an open-bid, ascending price auction akin to an English auction [14], with a 12-second deadline. For each auction instance, builders continuously submit bids in the form of $(B,BV)$ specifying a block $B$ and a bid value (BV) the builder is willing to pay for $B$ to be added to the blockchain. The revenue a bidder gains from $B$ is called the true value ($TV(B)$), which we estimate from transactions in $B$. At the end of an auction, the block $B$ with the highest bid is added to the blockchain, the builder pays $BV$ to the proposer and earns $TV$ when $B$ is executed (i.e., the builder’s profit is $TV-BV$).
@@
The two types of losses defined above can be computed from auction bids, as illustrated in Fig. 2. Specifically, suppose the true values of the bids from a given auction instance are $TV_{1}>TV_{2}\geq\cdots\geq TV_{N}$. The loss from inequality is $TV_{2}-TV_{1}$ because a competitive auction can, at most, generate revenue of $TV_{2}$. The loss from uncompetitiveness is $\max(TV_{2}-BV,0)$; if $TV_{2}-BV$ is negative, the auction is considered competitive without this loss.
@@
1.2 Private Order Flows
@@
In the current MEV ecosystem (c.f. Sec. 2.2), builders receive order flows from order flow providers and compute the optimal ordering of received order flows (see [17] for a reference builder algorithm). Thus, a builder’s block-building capacity is determined by the “quality” of the order flows it receives. At the same time, not all order flows are accessible to all builders. The difference in the builder’s ability to access important private order flows can drive the inequality in block-building capacity.
Not every provider is equally important, and some providers who generate profitable order flows can significantly impact builders. Thus, we propose a metric to identify important providers and investigate their implications.
Identifying pivotal private order flow providers.   To formally capture the intuition of “important providers”, we defined pivotal level to quantify the sustained impact of a given provider. A provider $P$ is pivotal in an auction $A$ if the winner of $A$ would have lost, had transactions from $P$ were not available. In other words, pivotal providers are necessary for the winner to win. The pivotal level of a provider $P$ is the percentage of MEV-Boost auctions in which $P$ is pivotal. Other metrics, calculated based on each provider’s contribution to the builders’ blocks [18, 19], may include temporarily influential providers, whereas this metric identifies sustainedly influential providers, resulting in fewer false positives.
@@
Order flows from these pivotal providers have two levels of accessibility, which we introduce below.
Accessibility by multiple builders.   Among the top 5 pivotal providers, MEV-Share and MEV Blocker send their order flows to all builders they consider reputable. They do not cause inequality loss since the competition among these builders allows proposers to capture this value.
The reputability of a builder is accessed per their specific rules. While top builders usually pass the assessments, it is harder for new builders to enter the market.
Exclusive pivotal providers.   A more worrisome finding is that some providers send order flows almost exclusively to their preferred builders. Besides the well-known exclusive relationships (known as “integration” [20]), we also identified that three of the top five pivotal providers, previously not known to have integration, began integrating with top builders in 2024. Our finding shows that the integration of these three pivotal providers accounted for at least 45% of the total proposer losses in 2024.
@@
1.3 Mitigation
Our study reveals trust concerns and incentive issues within the MEV supply chain. They cannot be resolved by the proposed alternative PBS designs [21, 13, 22, 23], as these designs may retain the same underlying supply chain.
To fully address these problems, we identify two essential properties: a security guarantee protecting providers from malicious builder attacks and a game-theoretical guarantee disincentivizing provider integration. While solving these problems remains a grand interdisciplinary challenge requiring further study, identifying these essential properties lays a foundation for future solutions.
Summary
To the best of our knowledge, we are the first to use the rich auction data to analyze the decentralization of Ethereum’s builder market, while most previous works are limited to on-chain data. As a result, they cannot identify the primary cause of proposer losses or accurately identify important providers. We summarize our contributions below.
  • We found that proposers incur significant losses in the centralized builder market, which could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles. We challenge the community’s belief that “builder centralization is okay”.
@@ 1.3 Mitigation
  • We identify a previously unrecognized integration of top providers, which causes inequality in block-building capacity and significant losses for proposers.
  • Our dataset of the true values of 301,479 MEV-Boost auctions from April to December 2023 may be of independent interest. Our dataset of 11 billion partial bids since 2022 is the largest scale known to us.
  • Besides contributing to the literature on MEV mitigation, our work also contributes to the Economics literature on the analysis of MEV-Boost auctions. The findings in this paper can serve as empirical evidence for future theoretical interrogation.
@@ 2.1 MEV-Boost Auctions
PoS Ethereum.   Before introducing background on MEV-Boost auctions, we first cover the relevant concepts in Ethereum’s Proof of Stake-based consensus protocol called Gasper [27]. Participants of Gasper are called validators. Each validator must deposit 32 ETH as collateral [28]. Time is divided into 12-second intervals called slots. This results in 7,200 slots in Ethereum per day. During each slot, a single validator, referred to as the proposer, proposes a block, while others vote for the chain head in their local view.
@@ 2.2 MEV Supply Chain
Channel.   A private channel, or channel for short, connects users directly with builders, bypassing the mempool. Channels receive user transactions and forward them to the user-specified builders, offering privacy and atomicity guarantees. E.g., a user can send a bundle (list) of transactions and request them not to be unbundled. Examples include MEV-Share [35], MEV Blocker [36], BloXroute [37], etc.
Searcher.   Searchers are a special type of users. They run (typically proprietary) algorithms to construct profitable transaction sequences (bundles) based on the pending transactions and the blockchain state. Same as regular users, searchers may submit transactions to mempool or channels. In addition, most builders have public APIs that allow searchers to submit bundles directly. Some searchers may prefer channels for the privacy and atomicity benefits. A searcher is said to integrate with a builder if it sends almost all of its orders to that builder exclusively. For example, SCP (addresses are listed in Tab. IV) is considered an integrated searcher of the builder Beaver [38].
@@
Note that $EV(B)$ is an estimation of the builder’s actual income from $B$ because the builder may own accounts out of $A_{p}$ in Equation 1, or even receive payments off-chain. We discuss the implications in Sec. 6.
@@ 3 Dataset
We created two auction datasets to study the decentralization of the builder market: a dataset of partial bids from September 2022 to October 2024, and a dataset of full bids from April to December 2023. The full bids are obtained from the ultra sound relay [15]. Recall from Sec. 2.1 that a bid consists of a block $B$ and a bid value $BV$ and a partial bid consists of $(\mathsf{Hash}(B),BV,PK_{\text{builder}},metadata)$ where $metadata$ includes parent hash, gas limit, gas used, etc. To support analysis, we also collected on-chain data, including blocks and transactions on Ethereum, as well as miscellaneous data, which will be detailed below. Tab. I summarizes the datasets.
@@
On-chain blocks 5,551,675 Reth node
transactions 852,482,466
Auction blocks produced by MEV-Boost 4,909,683 relay API
partial bids 11,966,097,921 relay API
full bids from ultra sound 326,574,676 ultra sound
Misc. builder public keys 273 see Sec. 3.1
@@
searcher addresses 2,172
private transactions 112,218,755
transaction sources 74,922,499
TABLE I: Overview of collected datasets.
@@ 3.1 Data Collection and Validation
On-chain data.   We run an Ethereum execution client Reth [47] along with a consensus client Lighthouse [48] to collect blocks and transactions from September 15, 2022, to October 31, 2024. The on-chain dataset includes 5,551,675 blocks and 852,482,466 transactions.
@@
Data from relay APIs.   We connected to all 13 known relays [^2] Three relays were inactive during the period of our study. and collected the following information from their public APIs: 1) the winning bid including the final block sent to the proposer; and 2) partial bids for blocks that did not win the auction. Recall that partial bids do not include the block body. We collected 4,909,683 winning bids and 11,966,097,921 partial bids from September 15, 2022, to October 31, 2024. We use winning bids to identify blocks produced through MEV-Boost and partial bids to quantify the importance of private order flow providers in Sec. 5.
@@
To validate the completeness of the data collected from relay APIs, we compared the winning bids we collected with three public datasets [49, 29, 50]. Our dataset includes 720 more winning bids than other datasets. Upon comparison with on-chain blocks, we find they are all included on Ethereum, indicating a high level of completeness.
@@
Full bids from ultra sound relay .   From ultra sound relay [15], we obtained the full bids in 301,479 historical auctions that took place between April and December 2023 (ultra sound relay stopped collecting bids at the end of 2023). Since the data volume is large (we record roughly 200 GB of data per day), our dataset samples one week from each month: April 9-15 and May 1-7 to December 1-7, 2023. The dataset comprises 326,574,676 full bids (block, bid value, and builder public key). We use full bids in Sec. 4 to quantify the competitiveness and efficiency of MEV-Boost auctions, as well as the proposers’ losses.
@@
The original data from relays includes bids submitted after the deadline due to network latency. We removed them as we are only interested in the bids before the auction ends. On April 18, 2023, MEV-Boost introduced bid cancellation that allows a bidder to cancel a previous bid by submitting a lower one. However, proposers can query bids before they are canceled and store them locally [51], effectively bypassing cancellation. Therefore we do not consider cancellation.
@@
Not all builders connect to the ultra sound relay, so we need to validate that the full bids from ultra sound relay cover a significant subset of all builders. We compute the percentage of the builders that submit to ultra sound relay over all known builders. The set of all known builders is a union of all builders that appeared in our datasets, including winning bids, partial bids, and bids from the ultra sound relay. As shown in Fig. 10, bids from the ultra sound relay covered more than 80% of the builders in over 75% of the MEV-Boost auctions from April 2023 to December 2023. Moreover, the top-5 builders by market share in this period (Flashbots, Beaver, Rsync, Builder0x69, and Titan) almost always submitted to ultra sound relay.
@@
  • Builder public keys: We collect public keys controlled by each builder combining two data sources: the public keys in builders’ official documents; and the distinct marks in their blocks’ “extra field” (see Appx. A for details).
    To validate the completeness of our dataset, we manually cross-check it with public datasets [52, 53] and related works [43, 38], and the results show that our dataset includes 154 more public keys, so our dataset is the most complete one so far.
@@
  • Searcher addresses: We obtained known searcher addresses from Etherscan [55], libMEV [56], and related works [38, 57].
@@
  • Transaction sources: we are interested in where a transaction originates from and which part of the MEV supply chain it has traversed through. We identified transaction sources using ground truth data (as described in Appx. A).
4 Proposer Loss in the MEV-Boost Auctions
In this section, we focus on the status quo and study the implications of a centralized builder market as we have today. In particular, we are interested in the losses incurred by proposers in the MEV-Boost auctions, as significant losses could lead to undesirable consequences, including validator centralization and inaccurate MEV oracles.
In principle, builder centralization does not imply a significant proposer loss, as optimal outcomes require only a few non-colluding builders with similar block-building capacities to maintain competitive auctions. To evaluate these two conditions in the centralized builder market, we quantify proposers’ losses in historical MEV-Boost auctions by considering two factors: the competitiveness of the auctions and the inequality in block-building capacity.